Accelerator Update for December 14, 2007 - December 17, 2007
Friday December 14
The day shift began with Operations monitoring the Tevatron (TeV) store 5800 (53.6E30), the Recycler stash (268.6E10), with the Antiproton Source (Pbar) stacking (30.3E10), with the Main Injector (MI), Booster, and the Linac operating normally, with the Collider Detector at Fermilab (CDF) and D-Zero (D0) using the colliding beams, and with NuMI , SciBooNE / MiniBooNE, and the Meson MTest experiments taking beam.
Operations terminated store 5800 at 10:42 AM.
Operations put the TeV into shot setup at 10:53 AM.
An RF expert reported that a Booster RF station (BRF15) was down due to a possible power amplifier problem and that BRF19 was turned on in its place.
At 12:27 PM, BRF10 tripped off for the first of four times in 11 minutes. After the last reset it was bypassed and an expert was asked to investigate.
Operations established store 5801 at 2:05 PM with an initial luminosity of 183.88E30.
A MI expert began a multi-batch slipstacking study at 3:48 PM.
Operations reported at 4 PM that the Recycler had stashed antiprotons twice during the day shift.
MI studies ended around 7 PM.
Operations reported at 11:59 PM that the Recycler had stashed three times during the evening shift.
Saturday December 15
The midnight shift began with Operations monitoring store 5801 (65.78E30) and the stash (225.3E10), with Pbar stacking (9.1E10), and with NuMI, Sci /MiniBooNE, and MTest taking beam.
Operations disabled the MTest beam at 12:05 AM.
Operations terminated store 5801 at 7:33 AM.
At 7:35 AM, Operations reported that no beam was making it into the Pbar Accumulator. A Pbar Debuncher injection septum (D:ISEPV) tripped off on an open interlock indication. Operators investigated and found a problem with a power supply. Operations contacted an expert.
Operations switched to a NuMI only timeline at 7:38 AM.
Operations put the TeV into shot setup at 7:47 AM.
Operations reported at 8:02 AM that the Recycler had stashed five times during the midnight shift.
Operations established store 5802 at 9:43 AM with an initial luminosity of 210.5E30.
At 10:55 AM, a Booster magnet (B:BEX3) tripped off on a Klixon indication. It would not reset. Operations contacted an expert.
A Pbar power supply expert brought D:ISEPV online at 11:47 AM.
At 1:20 PM, Booster experts reported that the Klixon trips were real, but no access was needed. These are air-cooled magnets and they are running at higher currents than in the past. Operators turned the magnet back on.
At 2:23 PM, the Booster East Anode Power Supply tripped off. It reset.
Operations disabled the MTest beam at 6:15 PM.
Operations reported at 7:59 PM that the Recycler had stashed four times during the last 12 hours.
At 10:52 PM, BEX3 tripped off. Operators let the magnet cool off for 30 minutes and turned it back on.
Sunday December 16
The midnight shift began with Operations monitoring store 5802 and a stash, with Pbar stacking, and with NuMI and Sci/MiniBooNE taking beam.
Operations established beam to MTest at 6:48 AM.
Operations terminated store 5802 at 7:29 AM.
Operations put the TeV into shot setup at 7:48 AM.
Operations reported at 8:07 AM that the Recycler had stashed six times during the last 12 hours.
Operations put the TeV through a dry squeeze at 10:04 AM.
Operations resumed sending beam to MTest at 10:30 AM.
At 1:34 PM, Operations reported that all the helium compressors at TeV sector A0 tripped off. Operators and the Duty Electrician investigated and discovered that the A0 pond water pump had a sheared shaft. Experts began searching for a spare pump.
Kicker experts reported at 2:41 PM that they had replaced a trigger module. The kicker was up and pulsing.
Operations confirmed with the Cryo System Coordinator that they could go back into shot setup without the A0 compressors. Operations put the TeV back into shot setup at 3:55 PM.
FESS technicians reported at 5:12 PM that they didn’t have a spare pump to replace the A0 pond water pump. As a temporary measure, they ran a fire hose from a fire hydrant into the A0 building to the helium compressors.
Operations established store 5805 at 6:06 PM with an initial luminosity of 211.0E30.
MTest experimenters reported at 7:53 PM that they were through with the beam and through with their test run.
Operations reported at 7:56 PM that the Recycler had stashed two times during the last 12 hours.
At 10:05 PM, a Booster power supply (B:MH2) tripped off for the first of many times over the rest of the shift. It reset each time.
Monday December 17
The midnight shift began with Operations monitoring store 5805 and a stash, with Pbar stacking, with B:MH2 continuing to trip off every 8 to 10 minutes, and with NuMI and Sci/MiniBooNE taking beam.
Cryo system experts reported at 12:02 AM that they had completed the A3 wet engine repair.
Operations put the TeV through a wet squeeze at 2:34 AM.
Operations put the TeV into shot setup at 4 AM.
Operations reported around 5 AM that B:MH2 stopped tripping. Why it was tripping off was never discovered.
At 5:23 AM, Operations reported that the Pelletron tripped off. Operators contacted an expert and with his help got the Electron Cooling operating.
Operations established store 5808 at 5:46 AM with an initial luminosity of 169.7E30.
Operations reported at 7:51 AM that the Recycler had stashed six times during the last hours.
The Plans for Monday
The plans for today are to stack, stash, and store. The Run Coordinator will let Booster, MI, and Swtichyard experts conduct studies.
Comments and Suggestions
|last modified 12/17/2007 email Fermilab|
|Security, Privacy, Legal|